if you have 3 chairs and 5 people how many different arrangements can you come up with?

Answer by jim_thompson5910(35256)

There are 5*4*3 = 20*3 = 60 different arrangements assuming that order matters. If order doesn't matter, then there are 60/(3*2*1) = 60/6 = 10 different arrangements.